OpenStudy (anonymous):

Express in terms of i. 4sqrt-225

OpenStudy (asnaseer):

do you know what 'i' represents?

OpenStudy (anonymous):

no the problem just stated express in in terms of i.

OpenStudy (asnaseer):

have you studied complex numbers?

OpenStudy (anonymous):

not that I remember

OpenStudy (asnaseer):

you need to understand them in order to answer this question.

OpenStudy (asnaseer):

have a look at this site to get an idea of what they are: http://www.mathsisfun.com/numbers/complex-numbers.html If you are still stuck after this then let me know which part is confusing and I will try and help you out.

OpenStudy (anonymous):

I still confused by the 4 at the begining of the equation

OpenStudy (anonymous):

would it be stated like 4i15

OpenStudy (asnaseer):

that just means multiply the final answer by 4.\[4\sqrt{-225}\]

OpenStudy (asnaseer):

so yes you got the square root correct \(\sqrt{-225}=-15i\), but you now need to multiply this by 4

OpenStudy (asnaseer):

sorry \(\sqrt{-225}=15i\)

OpenStudy (anonymous):

ok so if the sqrt of -225 is 15i then I multiply that by four so then the answer would be 60i

OpenStudy (asnaseer):

perfect! well done. :)

OpenStudy (anonymous):

ok thank you

OpenStudy (asnaseer):

yw :)